# Mesquite MoCap: Democratizing Real-Time Motion Capture with Affordable, Bodyworn IoT Sensors and WebXR SLAM

### v1

@@ I Introduction
Motion capture (MoCap) technology enables the recording, replay, and analysis of human movement, with applications spanning biomechanical research, clinical rehabilitation, animation, and interactive entertainment [kitagawa2020]. Traditional optical motion capture systems provide high-precision data but require dedicated studio spaces, optimal lighting conditions, and investments often exceeding \$100,000 [meng2021]. While IMU-based commercial alternatives offer portability, they remain costly due to proprietary hardware and software ecosystems [riener2016].
These constraints create significant accessibility barriers for many potential users, including small research laboratories, educational institutions, healthcare providers, and independent developers. Users seeking affordable alternatives often sacrifice customization and specificity, forcing them to compromise on both affordability and functionality. The limitations of existing motion capture systems stem from their reliance on expensive proprietary hardware and software, which restricts adaptability [cheung2016]. Furthermore, the complexity of setup, calibration, and the need for dedicated spaces can deter users from fully utilizing the technology. As a result, many researchers and practitioners are left without effective tools for capturing and analyzing human movement in real-world environments.
The Internet of Things (IoT) paradigm [atzori2010], along with Body Worn Sensor Networks (BWSNs) [pantelopoulos2010survey], offers a promising approach to address these limitations through distributed sensing, edge computing, and wireless data aggregation. By leveraging advances in microelectromechanical systems (MEMS) [foxlin2005], low-power wireless microcontrollers, and emerging web standards (such as WebSerial [webserial], WebSockets [websockets], WebXR [webxr], WebGL [webgl], and Progressive Web Apps [pwa]), it becomes possible to create affordable motion capture systems without compromising performance.
@@ II-A IMU-Based Motion Capture
Inertial Measurement Unit (IMU)-based motion capture has gained significant attention as a flexible and cost-effective alternative to traditional optical systems, which typically require complex setups and controlled environments. Unlike optical systems which rely on cameras and markers, IMU-based systems offer greater mobility, reduced infrastructure requirements, and the potential for use in unconstrained environments [filippeschi2017]. Commercial solutions like Xsens MVN [roetenberg2019], Perception Neuron [perception2021], and Rokoko Smartsuit Pro [rokoko] have demonstrated the viability of these approaches for full-body motion capture in entertainment, sports, and rehabilitation contexts. However, their high costs remain a barrier to broader adoption, particularly in educational, hobbyist, or low-resource research settings.
@@
To address this, several research efforts have explored low-cost alternatives. Smartphone-based systems [gowing2014, chen2018smartphone] and custom-built platforms using wireless microcontrollers [tang2015, liu2020] aim to lower the cost of entry but often trade off extensibility, ease and mature toolchains that integrate with existing software. Mesquite addresses these issues by exporting both raw data and the industry standard Biovision Hierarchy(BVH) format[meredith2001motion] which allows for easy integration with existing animation and motion capture software. Recent advancements in MEMS (Micro-Electro-Mechanical Systems) technology have improved the sensitivity, size, and cost of consumer-grade IMUs, making them suitable for precise motion tracking when combined with effective sensor fusion algorithms[madgwick2011, foxlin2005]. These algorithms enable real-time orientation estimation by fusing data from accelerometers, gyroscopes, and magnetometers. These operations are increasingly performed on the sensor node itself, reducing the need for high-bandwidth wireless communication by using on-board motion processing[roetenberg2019].
@@
Despite these improvements, existing low-cost solutions continue to face critical limitations. These include drift accumulation over time due to integration errors, restricted wireless communication architectures that hinder multi-sensor synchronization, and a general lack of robust, openly available software and hardware platforms that can be customized for application-specific needs [filippeschi2017, pfeil2016]. Moreover, few systems offer comprehensive real-time visualization or support real time streaming out-of-the-box.
@@
The core problem of IMU based mocap is calculating the position and orientation of the user in 3D space. While IMUs can provide accurate local motion data, they often struggle with drift and cumulative errors over time, making it difficult to maintain an accurate global reference frame. This is particularly challenging in dynamic environments where the user may move through complex trajectories or interact with physical objects. To solve this Mesquite can use any of the many implementations of Simultaneous Localization and Mapping (SLAM) techniques, which use computer vision to establish a global reference frame based on the user’s environment. This approach allows for accurate tracking of the user’s position and orientation in 3D space, even in dynamic environments. The system uses WebXR World Mapping [webxr, webxr-wm] to establish an absolute spatial frame, which is then used to track the user’s hip position. This approach allows for accurate tracking of the user’s position in 3D space, even in dynamic environments, while the rest of the body is tracked using 15 body-worn IMUs for orientation, and joint positions are estimated through forward kinematics.
@@ II-B IoT Sensor Networks for Human Motion Analysis
The Internet of Things (IoT) paradigm has catalyzed innovative approaches to sensor deployment and distributed data collection for human motion analysis. In particular, Body Sensor Networks (BSNs)—wireless networks of wearable sensors—have been extensively developed for a range of applications including healthcare monitoring [pantelopoulos2010survey], physical rehabilitation [yang2014], activity recognition [lara2013survey], and sports performance analysis [baca2010sensor]. These systems typically consist of multiple miniaturized wireless sensor nodes attached to various parts of the body, communicating with a central aggregator or hub that performs data fusion and higher-level analysis.
Several wireless communication protocols have been employed in BSNs, including Bluetooth Classic, Bluetooth Low Energy (BLE), ZigBee, Wi-Fi, and proprietary RF solutions [otanes2019]. Each protocol involves trade-offs between power consumption, data rate, communication range, and device complexity. For example, BLE is widely used due to its low power profile, but it may not support the high data throughput from multiple nodes required for high-frequency motion capture involving multiple sensors [milenkovic2006wireless].
A major challenge in motion capture applications is the requirement for high sampling rates (typically 30–120 Hz), low-latency data transmission, and tight synchronization across multiple sensors. Many existing BSNs are optimized for long battery life rather than real-time performance, which limits their suitability for dynamic motion capture tasks such as full-body tracking in eXtended Reality (XR) and biomechanical analysis [park2010wireless].
To address these limitations, edge computing approaches have been proposed, in which preliminary data processing—such as sensor fusion, noise filtering, or compression—is performed on the sensor node or a nearby edge device before transmitting results to the central system. This strategy can reduce wireless bandwidth requirements, mitigate latency, and improve the system’s overall responsiveness [rahmani2018edge]. Platforms such as Fog computing and lightweight embedded AI models are beginning to bridge the gap between real-time responsiveness and low-power operation in distributed sensing environments [pham2021iot]. In our case, the Mesquite system leverages ICM20948 IMUs with integrated Digital Motion Processors (DMPs) to perform onboard sensor fusion, significantly reducing the amount of data transmitted over the network and enabling real-time performance in resource-constrained environments.
@@ II-C Position and Orientation Estimation
IMU-based tracking systems rely on sensor fusion algorithms that combine data from accelerometers, gyroscopes, and magnetometers to estimate orientation. Algorithms such as the Kalman filter and its variants, including the Extended Kalman Filter (EKF), are widely used to optimally estimate the system state from noisy measurements [welch2006]. Due to the limitations of Euler angles (e.g., gimbal lock), quaternion-based representations are typically employed for robust and continuous orientation tracking [shoemake1985].
Recent advances in sensor hardware, such as the BNO08x (Bosch) [boschbno] and ICM-20948 (TDK) [tdk20948], have made it feasible to perform onboard sensor fusion using integrated motion processors. This not only improves accuracy but also reduces wireless data transmission overhead, a critical factor for real-time applications and networked sensor systems [roetenberg2019, boschbno].
@@
Inspite of these advancements, position estimation remains a significantly harder problem than orientation. It typically requires double integration of accelerometer signals, which amplifies noise and leads to rapid drift over time [foxlin2005, sabatini2011]. In the absence of external spatial references (e.g., GPS, beacons, or vision-based tracking), pure inertial positioning becomes unreliable over even short durations.
To overcome this, the Mesquite system adopts a hybrid approach by anchoring the user’s hip position using WebXR World Mapping [webxr, webxr-wm], which leverages computer vision SLAM techniques (e.g., Visual Odometry, sensor fusion, on-board ML modeling) from a hip-mounted smartphone to establish an absolute spatial frame. The rest of the body bones are tracked using 15 body-worn IMUs for orientation, and joint positions are estimated through forward kinematics, reducing the need for full-body positional tracking. This design choice exploits the fact that for many applications, accurate tracking of the hip position is sufficient, as the rest of the skeleton can be reconstructed using only bone orientations [gonzalez2020]. In order to capture a full motion frame we only need the position/orientation of the hip and the orientation of the rest of the bones of the body. This approach allows for a more practical and cost-effective solution while maintaining high accuracy in motion capture. This data is exported in BVH format, which is widely used in animation and motion analysis software. The raw data of each sensor is also logged for further analysis and post-processing if needed.
@@ III-D Skeletal Model and Animation
The skeletal model is visualized in real-time using WebGL, with the browser interface rendering the 3D animation based on the processed data. To improve tracking accuracy and mitigate common IMU limitations, the system employs multiple filtering approaches. Kalman filtering is implemented on the ESP32 to reduce noise in orientation estimation and mitigate gyroscope drift [welch2006]. Spherical linear interpolation (SLERP) is applied for smooth transitions between quaternion states [shoemake1985]. The system does not require magnetometer data for regular operation, making it robust to magnetic interference commonly found in indoor environments. However, magnetometer readings can optionally be incorporated to prevent long-term heading drift in extended recording sessions.
@@ III-E Data Export and Interoperability
The system exports motion data in the Biovision Hierarchy (BVH) format, enabling interoperability with existing animation and analysis software. The export pipeline includes quaternion to Euler angle conversion for compatibility with standard BVH implementations, joint hierarchy definition matching common animation standards, frame data export at configurable frame rates (15-45 FPS), and metadata inclusion for raw sensor data. The exported BVH files are compatible with software including Blender[blender], Maya[maya], Unity[unity], Unreal Engine[unreal], and other specialized biomechanical analysis tools. The system also supports CSV export for raw data analysis, allowing users to access the underlying sensor data for custom processing or visualization. This flexibility enables researchers and developers to integrate Mesquite into their existing workflows and leverage the system’s capabilities for a wide range of applications.
@@ IV Evaluation
To validate the accuracy of Mesquite, we conducted comparative benchmarking tests against an OptiTrack optical motion capture system[optitrack], widely considered the gold standard in motion tracking technology. The benchmarking methodology included simultaneous capture of a variety of human movements ranging from simple actions (walking, reaching) to complex motions (climbing, dancing etc).

### v2

@@ I Introduction
Motion capture (MoCap) technology enables the recording, replay, and analysis of human movement, with applications spanning biomechanical research, clinical rehabilitation, animation, and interactive entertainment [17]. Traditional optical motion capture systems provide high-precision data but require dedicated studio spaces, optimal lighting conditions, and investments often exceeding \$100,000 [21]. While IMU-based commercial alternatives offer portability, they remain costly due to proprietary hardware and software ecosystems [33].
These constraints create significant accessibility barriers for many potential users, including small research laboratories, educational institutions, healthcare providers, and independent developers. Users seeking affordable alternatives often sacrifice customization and specificity, forcing them to compromise on both affordability and functionality. The limitations of existing motion capture systems stem from their reliance on expensive proprietary hardware and software, which restricts adaptability [7]. Furthermore, the complexity of setup, calibration, and the need for dedicated spaces can deter users from fully utilizing the technology. As a result, many researchers and practitioners are left without effective tools for capturing and analyzing human movement in real-world environments.
The Internet of Things (IoT) paradigm [1], along with Body Worn Sensor Networks (BWSNs) [28], offers a promising approach to address these limitations through distributed sensing, edge computing, and wireless data aggregation. By leveraging advances in microelectromechanical systems (MEMS) [10], low-power wireless microcontrollers, and emerging web standards (such as WebSerial [13], WebSockets [15], WebXR [42], WebGL [16], and Progressive Web Apps [12]), it becomes possible to create affordable motion capture systems without compromising performance.
@@ II-A IMU-Based Motion Capture
Inertial Measurement Unit (IMU)-based motion capture has gained significant attention as a flexible and cost-effective alternative to traditional optical systems, which typically require complex setups and controlled environments. Unlike optical systems which rely on cameras and markers, IMU-based systems offer greater mobility, reduced infrastructure requirements, and the potential for use in unconstrained environments [9]. Commercial solutions like Xsens MVN [34], Perception Neuron [26], and Rokoko Smartsuit Pro [35] have demonstrated the viability of these approaches for full-body motion capture in entertainment, sports, and rehabilitation contexts. However, their high costs remain a barrier to broader adoption, particularly in educational, hobbyist, or low-resource research settings.
@@
To address this, several research efforts have explored low-cost alternatives. Smartphone-based systems [14, 6] and custom-built platforms using wireless microcontrollers [38, 19] aim to lower the cost of entry but often trade off extensibility, ease and mature toolchains that integrate with existing software. Mesquite addresses these issues by exporting both raw data and the industry standard Biovision Hierarchy(BVH) format[22] which allows for easy integration with existing animation and motion capture software. Recent advancements in MEMS (Micro-Electro-Mechanical Systems) technology have improved the sensitivity, size, and cost of consumer-grade IMUs, making them suitable for precise motion tracking when combined with effective sensor fusion algorithms[20, 10]. These algorithms enable real-time orientation estimation by fusing data from accelerometers, gyroscopes, and magnetometers. These operations are increasingly performed on the sensor node itself, reducing the need for high-bandwidth wireless communication by using on-board motion processing[34].
@@
Despite these improvements, existing low-cost solutions continue to face critical limitations. These include drift accumulation over time due to integration errors, restricted wireless communication architectures that hinder multi-sensor synchronization, and a general lack of robust, openly available software and hardware platforms that can be customized for application-specific needs [9, 30]. Moreover, few systems offer comprehensive real-time visualization or support real time streaming out-of-the-box.
@@
The core problem of IMU based mocap is calculating the position and orientation of the user in 3D space. While IMUs can provide accurate local motion data, they often struggle with drift and cumulative errors over time, making it difficult to maintain an accurate global reference frame. This is particularly challenging in dynamic environments where the user may move through complex trajectories or interact with physical objects. To solve this Mesquite can use any of the many implementations of Simultaneous Localization and Mapping (SLAM) techniques, which use computer vision to establish a global reference frame based on the user’s environment. This approach allows for accurate tracking of the user’s position and orientation in 3D space, even in dynamic environments. The system uses WebXR World Mapping [42, 24] to establish an absolute spatial frame, which is then used to track the user’s hip position. This approach allows for accurate tracking of the user’s position in 3D space, even in dynamic environments, while the rest of the body is tracked using 15 body-worn IMUs for orientation, and joint positions are estimated through forward kinematics.
@@ II-B IoT Sensor Networks for Human Motion Analysis
The Internet of Things (IoT) paradigm has catalyzed innovative approaches to sensor deployment and distributed data collection for human motion analysis. In particular, Body Sensor Networks (BSNs)—wireless networks of wearable sensors—have been extensively developed for a range of applications including healthcare monitoring [28], physical rehabilitation [43], activity recognition [18], and sports performance analysis [3]. These systems typically consist of multiple miniaturized wireless sensor nodes attached to various parts of the body, communicating with a central aggregator or hub that performs data fusion and higher-level analysis.
Several wireless communication protocols have been employed in BSNs, including Bluetooth Classic, Bluetooth Low Energy (BLE), ZigBee, Wi-Fi, and proprietary RF solutions [27]. Each protocol involves trade-offs between power consumption, data rate, communication range, and device complexity. For example, BLE is widely used due to its low power profile, but it may not support the high data throughput from multiple nodes required for high-frequency motion capture involving multiple sensors [23].
A major challenge in motion capture applications is the requirement for high sampling rates (typically 30–120 Hz), low-latency data transmission, and tight synchronization across multiple sensors. Many existing BSNs are optimized for long battery life rather than real-time performance, which limits their suitability for dynamic motion capture tasks such as full-body tracking in eXtended Reality (XR) and biomechanical analysis [29].
To address these limitations, edge computing approaches have been proposed, in which preliminary data processing—such as sensor fusion, noise filtering, or compression—is performed on the sensor node or a nearby edge device before transmitting results to the central system. This strategy can reduce wireless bandwidth requirements, mitigate latency, and improve the system’s overall responsiveness [32]. Platforms such as Fog computing and lightweight embedded AI models are beginning to bridge the gap between real-time responsiveness and low-power operation in distributed sensing environments [31]. In our case, the Mesquite system leverages ICM20948 IMUs with integrated Digital Motion Processors (DMPs) to perform onboard sensor fusion, significantly reducing the amount of data transmitted over the network and enabling real-time performance in resource-constrained environments.
@@ II-C Position and Orientation Estimation
IMU-based tracking systems rely on sensor fusion algorithms that combine data from accelerometers, gyroscopes, and magnetometers to estimate orientation. Algorithms such as the Kalman filter and its variants, including the Extended Kalman Filter (EKF), are widely used to optimally estimate the system state from noisy measurements [41]. Due to the limitations of Euler angles (e.g., gimbal lock), quaternion-based representations are typically employed for robust and continuous orientation tracking [37].
Recent advances in sensor hardware, such as the BNO08x (Bosch) [5] and ICM-20948 (TDK) [39], have made it feasible to perform onboard sensor fusion using integrated motion processors. This not only improves accuracy but also reduces wireless data transmission overhead, a critical factor for real-time applications and networked sensor systems [34, 5].
@@
Inspite of these advancements, position estimation remains a significantly harder problem than orientation. It typically requires double integration of accelerometer signals, which amplifies noise and leads to rapid drift over time [10, 36]. In the absence of external spatial references (e.g., GPS, beacons, or vision-based tracking), pure inertial positioning becomes unreliable over even short durations.
To overcome this, the Mesquite system adopts a hybrid approach by anchoring the user’s hip position using WebXR World Mapping [42, 24], which leverages computer vision SLAM techniques (e.g., Visual Odometry, sensor fusion, on-board ML modeling) from a hip-mounted smartphone to establish an absolute spatial frame. The rest of the body bones are tracked using 15 body-worn IMUs for orientation, and joint positions are estimated through forward kinematics, reducing the need for full-body positional tracking. This design choice exploits the fact that for many applications, accurate tracking of the hip position is sufficient, as the rest of the skeleton can be reconstructed using only bone orientations [11]. In order to capture a full motion frame we only need the position/orientation of the hip and the orientation of the rest of the bones of the body. This approach allows for a more practical and cost-effective solution while maintaining high accuracy in motion capture. This data is exported in BVH format, which is widely used in animation and motion analysis software. The raw data of each sensor is also logged for further analysis and post-processing if needed.
@@ III-D Skeletal Model and Animation
The skeletal model is visualized in real-time using WebGL, with the browser interface rendering the 3D animation based on the processed data. To improve tracking accuracy and mitigate common IMU limitations, the system employs multiple filtering approaches. Kalman filtering is implemented on the ESP32 to reduce noise in orientation estimation and mitigate gyroscope drift [41]. Spherical linear interpolation (SLERP) is applied for smooth transitions between quaternion states [37]. The system does not require magnetometer data for regular operation, making it robust to magnetic interference commonly found in indoor environments. However, magnetometer readings can optionally be incorporated to prevent long-term heading drift in extended recording sessions.
@@ III-E Data Export and Interoperability
The system exports motion data in the Biovision Hierarchy (BVH) format, enabling interoperability with existing animation and analysis software. The export pipeline includes quaternion to Euler angle conversion for compatibility with standard BVH implementations, joint hierarchy definition matching common animation standards, frame data export at configurable frame rates (15-45 FPS), and metadata inclusion for raw sensor data. The exported BVH files are compatible with software including Blender[4], Maya[2], Unity[40], Unreal Engine[8], and other specialized biomechanical analysis tools. The system also supports CSV export for raw data analysis, allowing users to access the underlying sensor data for custom processing or visualization. This flexibility enables researchers and developers to integrate Mesquite into their existing workflows and leverage the system’s capabilities for a wide range of applications.
@@ IV Evaluation
To validate the accuracy of Mesquite, we conducted comparative benchmarking tests against an OptiTrack optical motion capture system[25], widely considered the gold standard in motion tracking technology. The benchmarking methodology included simultaneous capture of a variety of human movements ranging from simple actions (walking, reaching) to complex motions (climbing, dancing etc).